# Forecasting Influenza Hospitalizations Using a Bayesian Hierarchical Nonlinear Model with Discrepancy

### v1

@@ Abstract
The annual influenza outbreak leads to significant public health and economic burdens making it desirable to have prompt and accurate probabilistic forecasts of the disease spread. The United States Centers for Disease Control and Prevention (CDC) hosts annually a national flu forecasting competition which has led to the development of a variety of flu forecast modeling methods. For the first several years of the competition, the target to be forecast was weekly percentage of patients with an influenza-like illness (ILI), but in 2021 the target was changed to weekly hospitalization counts. Reliable state and national hospitalization data has only been available since 2021, but for ILI the data has been available since 2010 and has been successfully forecast for several seasons. In this manuscript, we introduce a two component modeling framework for forecasting weekly hospitalizations utilizing both hospitalization data and ILI data. The first component is for modeling ILI data using a dynamic nonlinear Bayesian hierarchical model. The second component is for modeling hospitalizations as a function of ILI. For hospitalization forecasts, ILI is first forecasted and then hospitalizations are forecast with ILI forecasts used as a linear or quadratic predictor. In a simulation study, two ILI forecast models, including one similar to the winning model for two seasons of the CDC forecast competition from Osthus et al. and a nonlinear Bayesian hierarchical model from Ulloa are compared. Also assessed is the usefulness of including a systematic model discrepancy term in the ILI model. Forecasts of state and national hospitalizations for the 2023-24 flu season are made, and different modeling decisions are compared. We found that including a discrepancy component in the ILI model tends to improve forecasts during certain weeks of the year. We also found that other modeling decisions such as the exact nonlinear function to be used in the ILI model or the error distribution for hospitalization models may or may not be better than other decisions, depending on the season, location, or week of the forecast.
@@ 1 Introduction
To better inform public decision making regarding the flu epidemic, in 2013 the United States Centers for Disease Control and Prevention (CDC) organized a national flu forecasting competition, also known as FluSight Biggerstaff et al. (2016); Mathis et al. (2024); CDC (2024a). Originally, over a dozen teams of researchers from academic and industry backgrounds participated in FluSight by contributing their own forecast models. Besides the 2020 season –or the flu season spanning the fall of 2020 and the winter of 2021– FluSight has been operated annually and researchers outside the CDC have been invited to participate. Initially the target data for forecasts was influenza-like illness (ILI) data. ILI is the proportion of patients who meet a healthcare provider and who display flu like symptoms, and ILI data has been available at the state and national level since the 2010 flu season CDC (2024c, b). The collaborative ILI forecasting effort has led to a number of modeling developments in flu forecasting (McAndrew and Reich, 2021; Osthus and Moran, 2021; Osthus et al., 2019; Ulloa, 2019, see references therein for more examples), and in their paper’s introduction, Osthus et al. Osthus et al. (2019) categorized the most commonly used flu forecasting models into four classes including mechanistic models based on differential equation compartmental models, agent based models based on population simulation, machine learning/regression models including data driven machine learning and statistical models, and data assimilation models which are constructed by assimilating mechanistic models into a probabilistic framework. An additional forecast model used in FluSight involves the combination of several forecasts into a single ensemble forecast, which has been shown to perform well relative to individual models McAndrew and Reich (2021); Ray et al. (2020); Yamana et al. (2016).
@@
The administration of FluSight saw few changes during the first seven seasons, but the onset of the COVID-19 pandemic and subsequent developments for COVID-19 forecasting led to major modifications. As a result of the COVID-19 pandemic which began during the 2019 flu season, the typical flu outbreak behavior was altered between the 2019 and 2022 seasons Mathis et al. (2024). The COVID-19 pandemic led to the creation of the Health and Human Services (HHS) Patient Impact and Hospital Capacity Data System HealthData.gov (2024) which contains COVID-19 and flu hospitalization data, and the COVID-19 Forecast Hub was founded. The COVID-19 Forecast Hub was based on FluSight but with certain major adjustments including how the forecast uncertainty is represented and the addition of the weekly publication of a multi-model ensemble forecast as the official forecast of the CDC Bracher et al. (2021); Cramer et al. (2022). Using estimated quantiles for representing forecast uncertainty and creating a multi-model ensemble are both aspects of the COVID-19 Forecast Hub which were adopted by the flu forecast competition. Additionally the target of the flu forecasts changed from being ILI data to being HHS hospitalization data, which reports the number of hospitalizations due to a laboratory confirmed flu infection Mathis et al. (2024); HealthData.gov (2024). This is as a result of having COVID-19 cases in the population making ILI data, already only a proxy for flu behavior, more difficult to interpret.
@@
The contribution of this manuscript is to introduce a two component framework for modeling HHS hospitalization forecasts where hospitalization data and years of ILI data are used to inform forecast models. The first modeling component is a model of ILI data and the second is a model of hospitalization data with ILI as a predictive covariate. Herein we use ILI models similar to those in Osthus et al. Osthus et al. (2019) and Ulloa Ulloa (2019) for ILI forecasting. The model of Osthus et al. Osthus et al. (2019) is a combined data assimilation and statistical regression model which involves a compartmental model in a probabilistic framework. Their model also includes an additional component for capturing a systematic discrepancy between the deterministic part of the model and the actual data, an idea which was first introduced by Kennedy and O’Hagan Kennedy and O’Hagan (2001). The model in Ulloa Ulloa (2019) is a Bayesian hierarchical regression model with an underlying function intended to capture the trajectory of the seasonal ILI data. Herein, we provide a framework under which discrepancy modeling may be used along with a general function modeling ILI data, and we show the effectiveness of including discrepancy modeling during certain periods of the flu season.
In line with the newer FluSight standard of forecasting hospitalizations, we model hospitalizations as a linear function of ILI. Thus forecasts produced herein target flu hospitalizations and are a mapping of ILI forecasts to hospitalizations. This allows for ILI data from many seasons to be exploited and for ILI forecasts to assist in forecasting hospitalizations, which has fewer seasons of data than ILI. Several modeling schemes and their forecasts for the 2023 flu hospitalization season are compared, and it is shown that the modeling decisions produce good forecast results for different states or times during the flu season.
@@
In section 2 we review the ILI and hospitalization data provided by the CDC and targeted by FluSight. In section 3 the modeling framework contributed by this manuscript is given. In the same section, functions similar to those used by Osthus et al. Osthus et al. (2019) and Ulloa Ulloa (2019) are defined. These functions are the susceptible-infectious-recovered (SIR) compartmental model and the asymmetric Gaussian (ASG) function respectively. Model fitting and implementation are described at the end of the section. Section 4 includes a simulation study where four ILI forecast models and their use in forecasting hospitalizations are compared. Commonly used proper scoring rules Gneiting and Raftery (2007), which are also introduced and defined in section 4, are used for comparing the forecasts. Forecasting of the 2023 flu outbreak along with assessment and comparison under several selected models is performed in section 5. Finally, the manuscript is concluded in section 6 with general observations and some discussion.
@@ 2 Flu outbreak data
In this section we introduce and define ILI and hospitalization data and visually evaluate the data. ILI and hospitalization data have been the object of forecasting for FluSight with ILI being the target for the first seven seasons and hospitalizations being the target since the 2022 season. Both of these data were collected at the state, territorial, and national level and were reported at least weekly. Overall the data is reported for 53 locations including the 50 US states, the District of Columbia (DC), Puerto Rico (PR), and at the US national level. We will refer to forecast targets throughout this manuscript. A target is the specific horizon, 1, 2, 3, or 4-weeks ahead, for a specific location and week during the season.
@@ 2.1 Influenza-like illness data
The US Outpatient Influenza-like Illness Surveillance Network (ILINet) collects information on respiratory illness from outpatient visits to health care providers. Over 3,400 outpatient health care providers in all 50 US states, PR, DC, and the US Virgin Islands report each week the total number of outpatient visits along with the number of ILI cases. An ILI case is defined as a “fever (temperature of $100^{\circ}$[$37.8^{\circ}$] or greater) and a cough and/or a sore throat.” Prior to the 2021 season, the definition included “without a known cause other than influenza” CDC (2024c). Because other illnesses such as COVID-19, RSV, and the common cold may induce similar respiratory symptoms, ILI may include patients infected with some disease other than influenza. To know whether or not a sick patient is infected with influenza requires a laboratory test.
In 2013, when FluSight began, the ILI data was the object of the forecasts. The data was released publicly at HHS region levels, and forecast teams were asked to provide forecasts of several ILI targets on the regional levels including season onset, 1-4 week ahead ILI levels, and the week of peak ILI activity Biggerstaff et al. (2016); McGowan et al. (2019). Currently, the ILI data is collected by the CDC and published on an online portal for viewing at the national, HHS region, census, and state levels CDC (2024b). To obtain ILI data, we used the R package cdcfluview which provides functions for downloading the data Rudis (2021). Weekly ILI data from the national, HHS region, and census levels are available from the 1997 flu season until the current season. At the state level, data is available from the 2010 flu season to the current season.
The top of figure 1 shows the ILI data at the national level for flu seasons 2010 to 2023. For most seasons there are 52 weeks listed, but for the 2010, 2015, and 2021 seasons there are 53 weeks because there were 53 Sundays during those seasons. To better align with the flu behavior, week 1 is set as the first week of August and week 52 or 53 is the last week in July of the following year. For example week 1 of the 2013 season corresponds to the first week of August 2013, and week 52 of the same season corresponds to the last week of July 2014. This convention is used for the remainder of this manuscript.
@@
Notable from the plots in the top of figure 1 is the regular trajectory of the ILI. With the exception of season 2020, the ILI begins low at week 1 and increases as the fall and winter progress until the ILI reaches a peak. As spring progresses to summer, the ILI decreases to low values. As Osthus et al. Osthus et al. (2019) point out, there is nearly always either a global or local peak at week 22 which typically corresponds to the week between Christmas and New Year’s day. Whether local or global, the ILI holiday peak is generally expected and thought to be due to widespread holiday travel, school closure, or other unique social behavior Ewing et al. (2017); Garza et al. (2013). The only seasons when there was not a peak at week 22 were season 2022 where the season peak occurred particularly early and season 2020 which was greatly influenced by the COVID-19 pandemic.
@@ 2.2 Hospitalization Data
Hospital admission data, used as the object of FluSight forecasting for the 2022 and 2023 seasons, is based on the CDC’s National Healthcare Safety Network (NHSN) dataset entitled HealthData.gov COVID-19 Reported Patient Impact and Hospital Capacity by State Timeseries. Several targets of respiratory illnesses including COVID-19, RSV, and influenza are reported weekly by most hospitals in the US. In February 2022 it became mandatory for all hospitals to report the number of COVID-19 and influenza hospitalizations, and since then reporting of hospitalizations has become widespread. These data were updated every Wednesday and Friday according to NHSN guidelines HealthData.gov (2024).
The bottom of figure 1 shows the weekly national hospitalizations for the 2022 and 2023 flu seasons. These plots show similarities to the ILI plots in the top of the figure in that at the early weeks of the season hospitalizations are low, but they increase in the fall to a peak after which they decrease until the flu outbreak ends. For both the 2022 and 2023 seasons, the hospitalizations peaked during the same week as ILI, and in 2023 that peak occurred during the holiday week 22. Additional plots of ILI data and hospitalization data for select states are included in the supplementary material Wadsworth and Niemi (2024). These plots show flu outbreak trajectories similar to those in figure 1 with activity increasing to a peak and then decreasing.
@@ 3 ILI and hospitalization forecast modeling
The typical behavior of the ILI data which starts at lower values in the late summer and fall but which increases to a peak, usually in December or January, followed by a decline motivates the use of a nonlinear function which follows a similar trajectory. Compartmental models are standard mathematical models used for modelling disease outbreaks. One important compartmental model is the susceptible-infectious-recovered (SIR) model, which is used by some to model ILI data Osthus et al. (2019); Allen (2017). Ulloa Ulloa (2019) chose to use the asymmetric Gaussian (ASG) function to model ILI data. The SIR and ASG models may both be appropriate to describe ILI behavior over the course of a flu season, however there may also be systematic behavior not captured by either, necessitating an additional model component to capture the discrepancy. In the first part of this section, we present an ILI model similar to the model in Osthus et al. Osthus et al. (2019). With some generalization, the model may incorporate any appropriate nonlinear function, though the focus here is on the SIR and ASG functions which are defined.
With the aim of forecasting hospitalizations, we also introduce a linear model of hospitalizations data with ILI as a predictive covariate. To forecast hospitalizations, ILI data is first forecasted and the forecast is then plugged in as a covariate in the hospitalization model, thus producing hospitalization forecasts. The hospitalization model is also defined in this section, and the section is concluded with descriptions of selected prior distributions, model implementation, and posterior sampling.

### v2

@@ Abstract
The annual influenza outbreak leads to significant public health and economic burdens making it desirable to have prompt and accurate probabilistic forecasts of the disease spread. The United States Centers for Disease Control and Prevention hosts annually a national flu forecasting competition which has led to the development of a variety of flu forecast modeling methods. For the first several years of the competition, the target to be forecast was weekly percentage of patients with an influenza-like illness (ILI), but in 2021 the target was changed to weekly hospitalization counts. Reliable state and national hospitalization data has only been available since 2021, but for ILI the data has been available since 2010 and has been successfully forecast for several seasons. In this manuscript, we introduce a two component Bayesian modeling framework for forecasting weekly hospitalizations utilizing both hospitalization data and ILI data. The first component is for modeling ILI data using a nonlinear Bayesian hierarchical model. This component includes modeling discrepancy between a nonlinear functional model and the observed ILI which requires thoughtful decisions for modeling constraints and selecting prior parameter distributions. The second component is for modeling hospitalizations as a function of ILI. For hospitalization forecasts, ILI is first forecasted and then hospitalizations forecasts are produced using ILI forecasts as a linear or quadratic predictor. The combination of the models is done in a manner similar to using the Bayesian cut. In a simulation study, two ILI forecast models, including one similar to the winning model for two seasons of the CDC forecast competition from Osthus et al. and a nonlinear Bayesian hierarchical model from Ulloa are compared with other standard forecasts models. The usefulness of including a systematic model discrepancy term in the ILI model is specifically addressed and proves to greatly improve forecasts. In a real data analysis, forecasts of state and national hospitalizations for the 2023-24 flu season are made and compared with forecasts from 20 other competing models. The forecasts from the proposed methodology outperformed all but one of 20 competing models.
Keywords Disease outbreak forecasting  $\cdot$ Bayesian hierarchical modeling  $\cdot$ Probabilistic forecasting  $\cdot$ Model discrepancy
@@ 1 Introduction
To better inform public decision making regarding the flu epidemic, in 2013 the United States Centers for Disease Control and Prevention (CDC) organized a national flu forecasting competition, also known as FluSight Biggerstaff et al. (2016); Mathis et al. (2024); CDC (2024b). The competition lasted from the fall of 2013 through spring of 2014, and originally over a dozen teams of researchers from academic and industry backgrounds participated in FluSight by contributing their own forecast models. Besides the 2020-21 season, FluSight has been operated annually and researchers outside the CDC have been invited to participate. From hereon, we denote a flu season by the year in which the outbreak began, i.e. the 2020 season season outbreak began in 2020 and lasted until spring 2021. Initially the target data for forecasts was influenza-like illness (ILI) data. ILI is the proportion of patients who meet a healthcare provider and who display flu like symptoms, and ILI data has been available at the state and national level since the 2010 flu season CDC (2024c, a). The collaborative ILI forecasting effort has led to a number of modeling developments in flu forecasting (McAndrew and Reich, 2021; Osthus and Moran, 2021; Osthus et al., 2019; Ulloa, 2019, see references therein for more examples), and in their introduction, Osthus et al. Osthus et al. (2019) categorized the most commonly used flu forecasting models into four classes including mechanistic models based on differential equation compartmental models, agent based models based on population simulation, machine learning/regression models including data driven machine learning and statistical models, and data assimilation models which are constructed by assimilating mechanistic models into a probabilistic framework. An additional forecast modeling class used in FluSight involves the combination of several forecasts into a single ensemble forecast, which has been shown to perform well relative to individual models McAndrew and Reich (2021); Ray et al. (2020); Yamana et al. (2016).
@@
The administration of FluSight saw few changes during the first seven seasons between 2013 and 2019, but the onset of the COVID-19 pandemic and subsequent developments for COVID-19 forecasting led to major modifications. As a result of the COVID-19 pandemic which began during the 2019 flu season, the typical flu outbreak behavior was altered between the 2019 and 2022 seasons Mathis et al. (2024). The COVID-19 pandemic led to the creation of the Health and Human Services (HHS) Patient Impact and Hospital Capacity Data System HealthData.gov (2024) which contains COVID-19 and flu hospitalization data, and the COVID-19 Forecast Hub was founded. The COVID-19 Forecast Hub was based on FluSight but with certain major adjustments including how the forecast uncertainty is represented and the addition of the weekly publication of a multi-model ensemble forecast as the official forecast of the CDC Bracher et al. (2021); Cramer et al. (2022). Using estimated quantiles for representing forecast uncertainty and creating a multi-model ensemble are both aspects of the COVID-19 Forecast Hub which were adopted by FluSight. Additionally the target of the flu forecasts changed from being ILI data to being HHS hospitalization data, which reports the number of hospitalizations due to a laboratory confirmed flu infection Mathis et al. (2024); HealthData.gov (2024). This in part was as a result of having COVID-19 cases in the population making ILI data, already only a proxy for flu behavior, more difficult to interpret.
@@
The contribution of this manuscript is to introduce a two component Bayesian modeling framework for modeling HHS hospitalization forecasts where hospitalization data and years of ILI data are used to inform forecast models. The first modeling component is a hierarchical model of ILI data. The second is a model of hospitalization data with ILI as a predictive covariate. Herein we use Bayesian ILI models similar to those in Osthus et al. Osthus et al. (2019) and Ulloa Ulloa (2019) for ILI forecasting. The model of Osthus et al. Osthus et al. (2019) is a combined data assimilation and statistical regression model which involves a compartmental model in a probabilistic framework. Their model also includes an additional component for capturing a systematic discrepancy between the deterministic part of the model and the actual data, an idea which was first introduced by Kennedy and O’Hagan Kennedy and O’Hagan (2001). The model in Ulloa Ulloa (2019) is a Bayesian hierarchical regression model with an underlying function intended to capture the trajectory of the seasonal ILI data. Herein, we provide a framework under which discrepancy modeling may be used along with a general function modeling ILI data. Modeling discrepancy introduced challenges which required thoughtful selection of model constraints and prior distributions to be used effectively.
In line with the newer FluSight standard of forecasting hospitalizations, we model hospitalizations as a linear function of ILI. The flu hospitalization forecasts are a linear mapping of ILI forecasts allowing for ILI data from many seasons to be exploited and assist in forecasting hospitalizations. To produce hospitalization forecasts, a scheme similar to the Bayesian cut is used where forecasts from the posterior predictive distribution of an ILI model are plugged into the hospitalization model to produce a posterior predictive hospitalization forecast distribution. We show that this scheme produces accurate and reasonably well calibrated probabilistic forecasts. The forecasts produced in this manuscript are compared with out of the box forecast models in a simulation study, and in a real data analysis the forecasts are compared with 20 other forecasts from the 2023 FluSight forecast competition. The proposed models outperform both the out of the box methods, and they outperform all but one of the FluSight models.
@@
In section 2 we review the ILI and hospitalization data provided by the CDC and targeted by FluSight. In section 3 the modeling framework contributed by this manuscript is given. In the same section, functions similar to those used by Osthus et al. Osthus et al. (2019) and Ulloa Ulloa (2019) are defined. These functions are the susceptible-infectious-recovered (SIR) compartmental model and the asymmetric Gaussian (ASG) function respectively. Model fitting and implementation are described at the end of the section. Section 4 includes a simulation study where four ILI forecast models and their use in forecasting hospitalizations are compared. Commonly used proper scoring rules Gneiting and Raftery (2007) are used for comparing the forecasts. Forecasting of the 2023 flu outbreak along with assessment and comparison of 20 FluSight models is performed in section 5. Finally, the manuscript is concluded in section 6 with general observations and some discussion.
@@ 2 Flu outbreak data
In this section we introduce, define, and visually evaluate ILI and hospitalization data. ILI and hospitalization data have been the object of forecasting for FluSight with ILI being the target for the first seven seasons, 2013 to 2019, and hospitalizations being the target since the 2022 season. Both of these data were collected at the state, territorial, and national levels and were reported at least weekly. Overall the data is reported for 53 locations including the 50 US states, the District of Columbia (DC), Puerto Rico (PR), and the US national level. We will refer to forecast targets throughout this manuscript. A target is the specific horizon, 1, 2, 3, or 4-weeks ahead, for a specific location and week during the season.
@@ 2.1 Influenza-like illness data
The US Outpatient Influenza-like Illness Surveillance Network collects information on respiratory illness from outpatient visits to health care providers. Each week, over 3,400 outpatient health care providers in all 50 US states, PR, DC, and the US Virgin Islands report the total number of outpatient visits along with the number of ILI cases. An ILI case is defined as a “fever (temperature of $100^{\circ}$F[$37.8^{\circ}$C] or greater) and a cough and/or a sore throat.“ Prior to the 2021 season, the definition included “without a known cause other than influenza“ CDC (2024c). Because other illnesses such as COVID-19, RSV, and the common cold may induce similar respiratory symptoms, ILI may include patients infected with some disease other than influenza. To know whether or not a sick patient is infected with influenza requires a laboratory test.
In 2013, when FluSight began, the ILI data was the object of the forecasts. The data was released publicly at HHS region levels, and forecast teams were asked to provide forecasts of several ILI targets on the regional levels including season onset, 1-4 week ahead ILI levels, and the week of peak ILI activity Biggerstaff et al. (2016); McGowan et al. (2019). Currently, the ILI data is collected by the CDC and published on an online portal for viewing at the national, HHS region, census, and state levels CDC (2024a). To obtain ILI data, we used the R package cdcfluview which provides functions for downloading the data Rudis (2021). Weekly ILI data from the national, HHS region, and census levels are available from the 1997 flu season up to the current season. At the state level, data is available from the 2010 flu season to the current season.
The top of figure 1 shows the ILI data at the national level for flu seasons 2010 to 2023. For most seasons there are 52 weeks listed, but for the 2010, 2015, and 2021 seasons there are 53 weeks because those seasons had 53 Sundays. To better align with the flu behavior, week 1 is set as the first week of August and week 52 or 53 is the last week in July of the following year. For example, week 1 of the 2013 season corresponds to the first week of August 2013, and week 52 of the same season corresponds to the last week of July 2014. This convention is used for the remainder of this manuscript.
@@
Notable from the plots in the top of figure 1 is the regular trajectory of the ILI. With the exception of season 2020, the ILI begins low at week 1 and increases as the fall and winter progress until the ILI reaches a peak. As spring progresses to summer, the ILI decreases to low values. As Osthus et al. Osthus et al. (2019) point out, there is nearly always either a global or local peak at week 22 which typically corresponds to the week between Christmas and New Year’s day. Whether local or global, the ILI holiday peak is generally expected and thought to be due to widespread holiday travel, school closure, or other unique social behavior Ewing et al. (2017); Garza et al. (2013). The only seasons when there was not a peak at week 22 were 2022, when the season peak occurred particularly early, and 2020 which was greatly influenced by the COVID-19 pandemic.
@@ 2.2 Hospitalization Data
Weekly hospital admission data, which was used as the object of FluSight forecasting for the 2022 and 2023 seasons, is based on the CDC’s National Healthcare Safety Network (NHSN) dataset entitled HealthData.gov COVID-19 Reported Patient Impact and Hospital Capacity by State Timeseries. In February 2022 it became mandatory for all hospitals to report the number of COVID-19 and influenza hospitalizations, and since then reporting of hospitalizations has become widespread. These data were updated every Wednesday and Friday according to NHSN guidelines HealthData.gov (2024). Weekly influenza hospitalization data is defined as the number of newly hospitalized patients with a confirmed diagnosis of influenza. From hereon, influenza hospitalizations will be referrred to as hospitalizations.
The bottom of figure 1 shows the weekly national hospitalizations for the 2022 and 2023 flu seasons. These plots show a trend similar to the ILI plots in that during the early weeks of the season hospitalizations are low, but they increase in the fall to a peak and then decrease until the flu outbreak ends. For both the 2022 and 2023 seasons, the hospitalizations peaked during the same week as ILI, and in 2023 that peak occurred during the holiday week 22. Additional plots of ILI data and hospitalization data for all 50 US states are included in the supplementary material Wadsworth and Niemi (2024). These plots show flu outbreak trajectories similar to those in figure 1.
@@ 3 ILI and hospitalization forecast modeling
The typical behavior of the ILI data to begin low, rise to a peak, and then fall motivates the use of a nonlinear function which follows a similar trajectory. Compartmental models are standard mathematical models used for modelling disease outbreaks which may capture ILI behavior. One important compartmental model is the susceptible-infectious-recovered (SIR) disease transmission model, which has been used by some to model ILI data Osthus et al. (2019); Allen (2017). The SIR is a mechanistic model used to capture transmission of a particular disease, and since ILI may be influenced by more diseases than influenza, the disease transmission described by the SIR model can only partially contribute to the true ILI data-generating mechanism Osthus et al. (2019).
Ulloa Ulloa (2019) chose to use the asymmetric Gaussian (ASG) function to model ILI data. The ASG is a smooth function which may approximate the same shape as the ILI trajectory. The ASG function has more parameters and is thus slightly more flexible than the SIR model, and thus may be able to roughly capture ILI. However, there may still be systematic behavior which it cannot capture. Thus when using either the SIR or ASG models, an additional component for modeling discrepancy should be included.
In the first part of this section, we present an ILI model similar to the model in Osthus et al. Osthus et al. (2019). With some generalization, the model may incorporate any appropriate linear or nonlinear function, though the focus here is on the SIR and ASG functions, defined later in this section. With the aim of forecasting hospitalizations, we also introduce a linear model of hospitalization data with ILI as a predictive covariate. To forecast hospitalizations, ILI data is first forecasted and the forecast is then, as a covariate, plugged into the hospitalization model to produce hospitalization forecasts.
Because we desire probabilistic forecasts, Bayesian modeling is a natural framework for describing forecast uncertainty. Though our main interest is in effeciently producing a probabilistic forecast and not necessarily in estimating model parameters, certain modeling decisions and prior distribution selections proved critical in producing good forecasts. After defining the ILI and hospitalization models, we describe the selection of prior distributions, model fitting, and producing forecasts from the posterior predictive distribution by combining ILI and hospitalization models using a scheme similar to the Bayesian cut Plummer (2015).